OpenStudy (anonymous):

DINGDONGLINGLONGPOTATO123CHINGCHANGCHONG

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ x}-\frac{ 1 }{ y} }{ \frac{ 1 }{ x }+\frac{ 1 }{ y } }\]

OpenStudy (anonymous):

It's easy. Take the LCM of the numerator and denominator. Both are xy.

OpenStudy (anonymous):

Alright.....:P how you do that?

OpenStudy (anonymous):

Then the expression becomes \[\frac{ \frac{ y-x }{ xy } }{ \frac{ y+x }{ xy } }\]

OpenStudy (anonymous):

Woah... can i eliminate the xy's?

OpenStudy (anonymous):

xy is both in numerator and denominator. It gets canceled.

OpenStudy (anonymous):

Yes, eliminate

OpenStudy (anonymous):

ok then. So my anwer is: y-x ---- y+x ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What answer do you need?

OpenStudy (anonymous):

that one is perfect i got it thanks :D

OpenStudy (anonymous):

You're welcome Gab.

OpenStudy (anonymous):

A MEDAL FOR JU.

OpenStudy (anonymous):

Yessssss.... It's my good day. :)

OpenStudy (anonymous):

:P i need some help on these questions cause im taking my final exam :P

OpenStudy (anonymous):

hahah.. okay! Ask. How many more?

OpenStudy (anonymous):

not sure im on question 23 of 53 but most of them are easy for me. :P

OpenStudy (anonymous):

Okay.. start asking

OpenStudy (anonymous):

What is the remainder when (2x^3 + 4x^2 – 32x – 40) ÷ (x – 4)

OpenStudy (anonymous):

This requires a bunch of synthetic division mumbo jumbo

OpenStudy (anonymous):

You just need answer?

OpenStudy (anonymous):

Sure i guess :D but explain the next one please

OpenStudy (anonymous):

WAIT.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lemme open another question so i can give you more deserved medals ;D

OpenStudy (anonymous):

mjb

OpenStudy (anonymous):

Remainder is 24

OpenStudy (anonymous):

:P okay :D A MEDAL FOR JU.